OpenStudy (anonymous):

PLEASE I NEED HELP ASAP. LIKE RIGHT NOW..... 1) A football stadium holds 52,000 fans. A college student is doing research and determines that on any given game day, the home team has five times as many fans as the visiting team. In order to help the student in his research, he represents the number of home team tickets as H and the visiting team’s tickets as V. Which system of equations does the college student use to determine how many tickets each team gets? a. V +5H = 52,000 b. V + 5H = 0 V +5H = 52,000 c. H + 5V = 52,000 d. H - 5V = 0 V +H = 52,000

OpenStudy (anonymous):

2) Find the minimum value of the function for the polygonal convex set determined by the given system of inequalities. 3x + 2y >= 14 -5x + 5y <= 10 -8x + 3y >= -29 f(x, y) = 8x + 8y a. The minimum is at (-5,5) b. The minimum is at (4,1) c. The minimum is at (3,2) d. The minimum is at (-5,2)